# BLaST: High Performance Inference and Pretraining using BLock Sparse Transformers

### v1

@@ 3.3. BSpMM Kernel
We now introduce our custom BSpMM kernel - the elementary algebraic building block that secures maximal speedup over dense baselines. The BSpMM kernel, while being part of our BLaST framework, can also be used independently as a stand-alone SpMM kernel, paving the road for fast, sparse linear algebra kernels across various domains. We then further extend it to the MLP blocks inside Transformer architectures to capture the most common compute pattern in ML workloads.
@@ 3.3.1. Kernel Outline
Figure 3. Overview of BLaST BSpMM. Our kernel employs a blocked Compressed Sparse Column (CSC) format, obtained by sparsifying neural network weights. BLaST leverages block-level, bottom-up 2D parallelism—implemented in Triton–to maximize GPU hardware resource utilization.
@@
Figure 3 shows the high-level design of our BLaST BSpMM kernel based on the SMaT framework (okanovic2024high, ). SMaT is an SpMM library that exploits Tensor Cores (TCs) for the efficient multiplication of unstructured sparse matrices. The block‑sparse implementation of SMaT relies on the low‑level CUDA Matrix Multiply‑Accumulate (MMA) API to extract maximum performance from GPUs. Its superior performance is attained by efficiently streaming thread‑level blocks, overlapping computation with data transfers, and fully engaging all available TC units through bottom‑up two‑dimensional parallelism (kwasniewski2019red, ) .
The main difference between BLaST BSpMM and SMaT BSpMM is that the sparsity structure is not imposed by the unstructured CSR input but rather arises from our custom sparsification strategy. As a result, we can bypass the key performance bottlenecks identified by Okanovic et al. (okanovic2024high, )—namely, the excessive fill-in caused by imposing an explicit block structure on CSR, and the highly skewed load imbalance among warps. The design of BLaST’s BSpMM kernel is guided by a rigorous performance model (okanovic2024high, ).
The carefully tuned rectangular block shape results from optimizing shared memory and register usage, while minimizing accuracy loss due to overly aggressive sparsification. The two block dimensions—blk_M, and blk_N—are optimized for compute performance, and model accuracy, respectively. Note that blk_N corresponds to block parameter $b$ introduced in Section 3.2.
@@
    blk_N determines the sparse block size. When the block is smaller, the sparsification strategy has more freedom to choose the best blocks to prune—resulting in a more fine-grained masking process. Conversely, larger blocks enable more efficient computation. We empirically select the largest block sizes that preserve high accuracy. This will be further explored in Section 5.4.1.
@@
    blk_M controls the number of rows from the dense matrix that participate in the threadblock-level reduction, thereby increasing the reuse of data loaded into shared memory from the sparse block. We use the COSMA analytical compute model (kwasniewski2019red, ) to derive the optimal block size, given hardware constraints such as the number of available registers, shared memory capacity, and warps per threadblock.
@@ 3.3.2. Implementation
The main challenge to implement a high-performance BCSC-based, batched SpMM kernel is to handle inherent data indirection efficiently. Compared to dense $Y=XW$, all memory accesses are known at the compile time, and we can create a provably optimal static schedule with compile time tilings, data distribution, and memory pipelining (ziogas2022deinsum, ). The sparse data structure prevents the static access pattern of the dense matrix $X$, since, to achieve data movement optimality, only necessary blocks of $X$ can be loaded - and the sparsity structure of $W$ determines this. We implement this in Triton by determining the dynamic range of accesses to $X$ with pointer algebra on the blk_col_ptr pointer array. The main part of the kernel is presented in Listing LABEL:lst:spmm-kernel. The Triton API can be found in its documentation (tillet2019triton, ).
@@ 3.3.3. Block Sparse MLP
In the MLP block, the linear layer is followed by a nonlinear operator. In most of machine learning, nonlinear operators fall into two categories – element-wise nonlinearities, such as ReLU, GeLU, SiLU, and row-wise operators, that perform some type of reduction, usually over the embedding dimension. Examples of such nonlinear operators are LayerNorm (Ba2016, ) or RMSNorm (Zhang2019, ).
While tensor contractions are usually compute-bound (even in the sparse case), nonlinear operators are memory-bound. The arithmetic intensity — the ratio of arithmetic operations per loaded element — is a small constant, usually 1 or 2. Therefore, to prevent a memory bottleneck, nonlinear operators have to be fused with the compute-bound operations.
Our optimized BSpMM kernel for MLPs fuses the nonlinearities that follow the SpMM operator into one kernel. Our implementation supports all common nonlinearities, further improving performance and reducing data movement.
@@ 4. Evaluation
We conduct a comprehensive set of experiments to evaluate the performance and accuracy of BLaST. Our experiments span 8 different datasets in both text and vision domain, across 9 different Transformer architectures.
In Section 5, we train the models on a diverse set of benchmark datasets, including OpenWebText (Gokaslan2019OpenWeb, ), CIFAR-10 (krizhevsky2009learning, ), LAMBADA (paperno2016lambadadatasetwordprediction, ), and five tasks from the GLUE benchmark (wang2019gluemultitaskbenchmarkanalysis, ). Furthermore, we conduct experiments using the following Transformer architectures along with their respective parameter counts: Llama3.2 (Panferov2025, ) (1.59B), GPT2-small (Radford2019, ) (124M), GPT2-medium (Radford2019, ) (355M), GPT2-large (Radford2019, ) (774M), GPT2-XL (Radford2019, ) (1.44B), ViT-B/16 (dosovitskiy2020image, ) (86M), ViT-B/32 (dosovitskiy2020image, ) (86M), ViT-L/16 (dosovitskiy2020image, ) (307M), ViT-L/32 (dosovitskiy2020image, ) (307M). In all our experiments, we adhere to the scientific benchmarking guidelines proposed by Hoefler et al. (hoefler2015scientific, ). That is, when reporting speedups we report the single parallel process as the base case.
Hardware Infrastructure.
We evaluate BLaST on a compute partition of the Alps supercomputer [^3] https://www.cscs.ch/computers/alps equipped with 2,688 compute nodes. Each node is equipped with four NVIDIA Grace Hopper GH200 superchips interconnected with HPE Cray Slinghot-11.
Software Stack.
All experiments were executed using Pytorch v2.7.0, CUDA v12.8, Megatron-core v0.10.0, Apex v0.1 and Transformers_engine v2.0.0.
@@ 5. Results
In Section 5.1, we present the speedups achieved by BLaST’s BSpMM compared to state-of-the-art sparse matrix-matrix multiplication algorithms, as well as the improvements observed within the MLP layers of LLMs scaled to 405B parameters. This section also showcases the end-to-end speedup achieved during LLM inference. Section 5.3 demonstrates the accuracies achieved during pretraining using BLaST’s Blocked and Prune algorithm. In Section 5.2, we show that BLaST can be used to recover the accuracy of previously pretrained models when inference speedup and memory reductions are of interest. Lastly, Section 5.4 presents an ablation study of BLaST, demonstrating its robustness to the choice of hyperparameters.
5.1. BLaST BSpMM Speedup
We begin by analyzing the speedup achieved by the BLaST BSpMM kernel relative to state-of-the-art matrix multiplication algorithms. Throughout this section, we report the speedup as the ratio
@@ 5.1. BLaST BSpMM Speedup
where $t_{CUTLASS}$ represents the execution time using the CUTLASS library, $t_{cuBLAS}$ denotes the time for cuBLAS, and $t_{BLaST{}}$ is the execution time for our method. In all experiments, we present the results for blk_M = blk_N $\in\{32,64,128\}$ which coincides to the TC MMA input sizes (see Section 3.3) as the sparsity level increases.
Figure 4 shows the speedup of the BSpMM kernel when multiplying individual matrices corresponding to typical embedding and context length sizes in LLMs. In a matrix multiplication routine characterized by the dimensions $M$, $N$, and $K$, the $Emb$ dimension corresponds to $K$, $Seq$ corresponds to $M$, and $N$ is set to $4\times Emb$. This indicates that as both the $Emb$ and $Seq$ dimensions increase, the speedup also increases. BLaST BSpMM is faster by as much as 16.7x for $Emb=32,768$, $Seq=1,024$, and sparsity 95%. We observe that, in general, higher performance is achieved with a larger block size; this plot uses a 128×128 block size. Moreover, we observe that BLaST SpMM with a 128×128 block size can even outperform cuBLAS by 1.1x in the fully dense case (0% sparsity). Note that BLaST BSpMM supports all data types supported by Triton (tillet2019triton, ). Due to the space constraints, we include the results only for the most popular data type in ML workloads: BF16. Using lower precision, e.g. torch.float8_e5m2, we measured up to 23x speedup for $Emb$ = 32,768, $Seq$ = 1,024, and sparsity 95%.
Lastly, we evaluate the speedup of BLaST BSpMM in the Llama model family, ranging from the Llama-3.2 1B to the Llama-3.1 405B. Figure 5 demonstrates the speedup in the MLP blocks. We observe that even 70% sparsity in the MLP blocks is sufficient to achieve a 2x speedup. As the models scale up to 405B parameters, we achieve up to an 8.8x speedup. In Figure 6, we show the end-to-end speedup in the inference of the Llama-3.2 1B model across different block sizes. One can achieve up to a 1.6x speedup even for a smaller model such as the Llama-3.2 1B using our proposed BSpMM. In conclusion, these experiments demonstrate that BLaST represents a promising step towards scaling models to the trillion-parameter era.
@@
Figure 6. Inference speedup for Llama-3.2 1B for various block sizes and sparsity levels. For 70% sparsity, BLaST yields 1.3x speedup, and for 95% sparsity, BLaST reaches up to 1.6x speedup.
@@ 5.2. BLaST for Inference
Pretraining neural networks from scratch demands substantial computational resources, especially for large dense models. Given the widespread availability of pretrained LLMs—which embody significant compute investments—there is a strong incentive to leverage these models rather than retrain a network from scratch. Our approach is capable of knowledge transfer by initializing the network with weights from an existing pretrained dense model. These pretrained weights may originate either from a dense network trained on the same dataset (the post-training compression setting) or from a network pretrained on a different dataset (the fine tuning setting). When training with pretrained dense weights, we have observed that, in cases involving large vocabulary sizes, utilizing a knowledge distillation loss improves the final accuracy (muralidharan2024compact, ; blakeman2025nemotron, ). That is, rather than using solely the cross-entropy loss $L_{CE}$, we backpropagate using a combined loss $\alpha L_{CE}+\beta L_{KL}$, where $L_{KL}$ represents the Kullback-Leibler divergence computed between the BLaST logits and those of the dense pretrained model, with weighting coefficients $\alpha,\beta\in\mathbb{R}$. In our experiments, we do not conduct an ablation study on the effect of knowledge distillation, as it is beyond the scope of this paper; instead, we report the best performance achieved between using, and not using the knowledge distillation loss.
In all experiments, we iteratively sparsify the model while simultaneously training it to recover the accuracy of the dense version (see Section 3). BLaST yields a model that maintains high accuracy while benefiting from faster inference and a reduced memory footprint. We demonstrate the capability of BLaST to recover the full accuracy of a pretrained dense network on downstream tasks.
In Table 1 we compare the performance of sparsifying BLaST under different levels of sparsity, and various block sizes on a subset of the GLUE benchmark.
BLaST CoLA SST-2 MRPC RTE WNLI Avg.
$s_{max}$ $b$ Matt. Corr ACC ACC / F1 ACC ACC Score
70% 32 41.59 84.98 72.06/80.94 59.57 56.34 65.91
80% 32 39.21 85.89 68.14/77.35 61.01 56.34 64.66
90% 32 27.78 84.63 70.34/79.17 61.37 56.34 63.27
95% 32 28.87 85.32 70.83/79.73 59.21 56.34 63.38
70% 64 34.22 84.40 66.42/77.28 59.21 56.34 62.98
80% 64 32.81 85.67 69.36/79.13 60.29 56.34 63.93
90% 64 24.01 84.29 68.87/77.36 58.12 56.34 61.50
95% 64 30.53 85.32 68.38/78.96 55.23 56.34 62.46
70% 128 36.70 86.24 70.83/79.86 60.65 56.34 65.10
80% 128 34.80 85.55 70.59/79.87 53.07 56.34 63.37
90% 128 34.76 85.09 69.12/80.25 59.21 56.34 64.13
95% 128 31.22 83.94 68.87/79.81 54.87 56.34 62.51
Dense 36.44 87.04 73.28/81.22 62.45 56.34 66.13
Table 1. Fine tuning Llama 3.2 1B with BLaST for GLUE benchmark. BLaST is robust to the choice of tile size, capable of lossless sparsification up to 95% for WNLI, and even outperforming Dense for CoLA.

### v2

@@ 3.4. BSpMM Kernel
We now introduce our custom BSpMM kernel, the elementary algebraic building block that secures maximal speedup over dense baselines. The BSpMM kernel, while being part of our BLaST framework, can also be used independently as a stand-alone SpMM kernel, paving the road for fast, sparse linear algebra kernels across various domains. We then further extend it to the MLP blocks inside Transformer architectures to capture the most common compute pattern in ML workloads.
@@ 3.4.1. Kernel Outline
Figure 4. Overview of BLaST BSpMM. The kernel uses a blocked Compressed Sparse Column (BCSC) layout derived from sparsified neural network weights. BLaST applies block-level, bottom-up 2D parallelism in Triton to maximize GPU utilization.
@@
Figure 4 illustrates the high-level design of our BLaST BSpMM kernel. We build on prior work such as SMaT (okanovic2024high, ), then focus on a block-sparse path tailored to our setting.
We store sparsified weights in a BCSC format. The launch grid is derived from BCSC so that we create work only for output tiles that have contributing nonzero blocks in $W$. Each output tile in $Y$ is assigned to one warp and driven by a single Tensor Core. We refer to this as bottom-up 2D parallelism because tiles exist only where $W$ has nonzero blocks. This reduces load imbalance and avoids work on empty regions.
For each output tile $Y_{i,j}$ our kernel calls low-level CUDA MMA (matrix multiply accumulate) instructions on the warp’s Tensor Core multiplying nonzero blocks $W_{:,j}$ with corresponding dense blocks $X_{i,:}$. The block size matches the MMA shape, enabling direct register-level fragments. We stream blocks with asynchronous copies into shared memory, and double buffer to overlap data movement with compute. Because iteration follows the BCSC column pointers, the kernel touches only existing blocks and never computes on absent tiles, which eliminates fill-in and keeps the Tensor Core pipelines busy.
The carefully tuned rectangular block shape results from optimizing shared memory and register usage, while minimizing accuracy loss due to overly aggressive sparsification. The two block dimension, blk_M, and blk_N, are optimized for compute performance, and model accuracy, respectively. Note that blk_N corresponds to block parameter $b$ introduced in Section 3.2:
@@
    blk_N determines the sparse block size. When the block is smaller, the sparsification strategy has more freedom to choose the best blocks to prune, resulting in a more fine-grained masking process. Conversely, larger blocks enable more efficient computation. We empirically select the largest block sizes that maintain accuracy, with a median relative accuracy drop of 2.24%.
@@
    blk_M controls the number of rows from the dense matrix that participate in the thread-block level reduction, thus increasing the reuse of data loaded into shared memory from the sparse block. We use the COSMA analytical compute model (kwasniewski2019red, ) to derive the optimal block size, given hardware constraints such as the number of available registers, shared memory capacity, and warps per threadblock.
@@ 3.4.2. Implementation
The main challenge to implement a high-performance BCSC-based, batched SpMM kernel is to handle inherent data indirection efficiently. Compared to dense $Y=XW$, all memory accesses are known at the compile time, and we can create a provably optimal static schedule with compile time tilings, data distribution, and memory pipelining (ziogas2022deinsum, ). The sparse data structure prevents the static access pattern of the dense matrix $X$, since, to achieve data movement optimality, only necessary blocks of $X$ can be loaded, and the sparsity structure of $W$ determines this. We implement this in Triton by determining the dynamic range of accesses to $X$ with pointer algebra on the blk_col_ptr pointer array. The main part of the kernel is presented in Listing LABEL:lst:spmm-kernel. The Triton API can be found in its documentation (tillet2019triton, ).
@@ 3.4.3. Block Sparse MLP
A Transformer MLP uses two affine projections with a element-wise nonlinearity in between, such as ReLU, GELU, or SiLU. While tensor contractions are usually compute-bound (even in the sparse case), nonlinear operators are memory-bound. The arithmetic intensity, the ratio of arithmetic operations per loaded element, is a small constant, usually 1 or 2. Therefore, to prevent a memory bottleneck, nonlinear operators have to be fused with the compute-bound operations.
Our optimized BSpMM kernel for MLPs fuses the nonlinearities that follow the SpMM operator into one kernel. Our implementation supports all common nonlinearities supported by Triton, further improving performance and reducing data movement.
@@ 4. Evaluation
We conduct a comprehensive set of experiments to evaluate the performance, and accuracy of BLaST. Our experiments span 8 different datasets in both text and vision domain, across 11 different Transformer architectures.
In Section 5, we train the models on a diverse set of benchmark datasets, including OpenWebText (Gokaslan2019OpenWeb, ), CIFAR-10 (krizhevsky2009learning, ), DialogSum (paperno2016lambadadatasetwordprediction, ), and five tasks from the GLUE benchmark (wang2019gluemultitaskbenchmarkanalysis, ).
Furthermore, we conduct experiments across both encoder and decoder-only Transformer architectures, along with their respective parameter counts: Llama3.2 (Panferov2025, ) (1.59B), Qwen3 (Yang2025a, ) (1.7B), BERT (Devlin2019, ) (110M), GPT2-small (Radford2019, ) (124M), GPT2-medium (Radford2019, ) (355M), GPT2-large (Radford2019, ) (774M), GPT2-XL (Radford2019, ) (1.44B), ViT-B/16 (dosovitskiy2020image, ) (86M), ViT-B/32 (dosovitskiy2020image, ) (86M), ViT-L/16 (dosovitskiy2020image, ) (307M), ViT-L/32 (dosovitskiy2020image, ) (307M). In all our experiments, we adhere to the scientific benchmarking guidelines proposed by Hoefler et al. (hoefler2015scientific, ). That is, when reporting speedups we report the single parallel process as the base case, and report the absolute performances of base cases in the appendix.
Hardware Infrastructure
We evaluate BLaST on a compute partition, where each node is equipped with four NVIDIA Grace Hopper GH200 superchips interconnected with HPE Cray Slinghot-11. Unless otherwise stated, measurements use a single GH200 superchip. Multi-node configurations, when reported, scale up to 4 nodes (16 GH200s).
Software Stack
All experiments were executed using Pytorch v2.7.0, CUDA v12.8, Megatron-core v0.10.0, Apex v0.1, Triton v3.0.0, and Transformers_engine v2.0.0.
Code and documentation are made available in the code repository: https://github.com/spcl/blast.
@@ 5. Results
5.1. BLaST BSpMM Kernel Comparison
Figure 5. Throughput of BLaST BSpMM compared to state-of-the-art kernels for various combinations of batch size, embedding and sequence length typically found in the networks that we benchmark. BLaST BSpMM can outperform all vendor implementations by a wide margin for all block sizes.
In this section, we benchmark the performance of our BLaST BSpMM kernel against state-of-the-art matrix multiplication kernels. We compare against PyTorch’s dense matrix multiplication implementation (paszke2019pytorch, ), which utilizes CUTLASS and cuBLAS libraries, PyTorch with Block Sparse Connectivity (BSC) sparsity format, as well as cuSPARSE (naumov2010cusparse, ) for sparse matrix operations. Figure 5 shows the throughput of BLaST (fixed 70%) for various matrix shapes over different matrix shapes, varying the embedding, sequence, and hidden dimensions as described in the previous chapter. This shows that BLaST BSpMM consistently outperforms all state-of-the-art kernels across the evaluated configurations. In particular, BLaST is capable of outperforming cuSPARSE by up to 521x, showcasing the significant computational advantages of our kernel.
5.2. BLaST BSpMM Performance
The previous section demonstrated the performance of our BSpMM kernel against other block SpMM kernels. Results show that dense kernels from cuBLAS and CUTLASS consistently rank second to our implementation. We now analyze the speedup achieved by the BLaST BSpMM kernel relative to state-of-the-art dense matrix multiplication algorithms, reporting speedup as the ratio
@@ 5.2. BLaST BSpMM Performance
where $t_{CUTLASS}$ represents the execution time using the CUTLASS library, $t_{cuBLAS}$ denotes the time for cuBLAS, and $t_{BLaST{}}$ is the execution time for our method. In all experiments, we present the results for blk_M = blk_N $\in\{16,32,64,128\}$.
Figure 6 shows the speedup of the BSpMM kernel when multiplying individual matrices corresponding to typical embedding and context length sizes in LLMs. In a matrix multiplication routine characterized by the dimensions $M$, NN, and $K$, the $Emb$ dimension corresponds to $K$, SeqSeq corresponds to $M$, and $N$ is set to $4\times Emb$. This indicates that as both the $Emb$ and $Seq$ dimensions increase, the speedup also increases. BLaST BSpMM is faster by as much as 16.7x for $Emb=32,768$, $Seq=1,024$, and sparsity 95%. We observe that, in general, higher performance is achieved with a larger block size. Moreover, we observe that BLaST SpMM with a 128×128 block size can even outperform cuBLAS by 1.1x in the fully dense case (0% sparsity). Note that BLaST BSpMM supports all data types supported by Triton (tillet2019triton, ). Due to the space constraints, we include the results only for the most popular data type in ML workloads: BF16. Using lower precision, e.g. torch.float8_e5m2, we measured up to 23x speedup for $Emb$ = 32,768, $Seq$ = 1,024, and sparsity 95%.
Lastly, we evaluate the speedup of BLaST BSpMM in the Llama model family, ranging from the Llama-3.2 1B to the Llama-3.1 405B. Figure 7 demonstrates the speedup in the MLP blocks. We observe that even 70% sparsity in the MLP blocks is sufficient to achieve a 2.15x speedup. As the models scale up to 405B parameters, we achieve up to an 8.77x speedup for 95% MLP sparsity. In Figure 8, we show the end-to-end speedup in the inference of the Llama-3.2 1B model across different block sizes. One can achieve up to a 1.6x speedup even for a smaller model such as the Llama-3.2 1B using our proposed BSpMM. In conclusion, these experiments demonstrate that BLaST represents a promising step towards scaling models to the trillion-parameter era.
@@
Figure 8. End-to-end inference speedup for Llama-3.2 1B for various block sizes and sparsity levels. For 70% sparsity, BLaST yields 1.3x speedup, and for 95% sparsity, BLaST reaches up to 1.6x speedup.
5.3. Distributed Results
In this section, we demonstrate the results of our method in a distributed setting. We benchmark the results on up to four nodes, where each node has four NVIDIA Grace Hopper GH200 superchips interconnected with HPE Cray Slinghot-11. BLaST can be adapted to all versions of parallelism used in deep learning (ben2019demystifying, ), since it is fundamentally a matrix multiplication method. For data parallelism, BLaST scales linearly with the number of GPUs as the model is replicated across each device, maintaining the same computational efficiency per GPU. The method naturally extends to tensor parallelism, where matrix operations are partitioned across devices, and model pipeline parallelism, where different layers are distributed across GPUs while preserving the sparsity patterns within each pipeline stage. Our method can be seamlessly integrated with any framework for distributed training. We evaluate our results using one of the most popular frameworks, Megatron-LM (shoeybi2019megatron, ), which provides efficient implementations of transformer models with support for various parallelization strategies including tensor, pipeline, and data parallelism. Figure 9 demonstrates our improvement over the dense baseline. We scaled up the Llama3 architecture according to the scaling laws (Kaplan2020, ), reaching 540B parameters for 16 GH200 NVIDIA GPUs. BLaST is 2.2x end-to-end faster compared to the dense model for 16 GPUs spread across 4 nodes for 95% sparsity in the MLP layers. Our results show that BLaST consistently outperforms the dense baseline across all distributed configurations.
Figure 9. Speedup of BLaST in the distributed setting using Megatron-LM. We scale the Llama 3 architecture up to 16 NVIDIA GH200 GPUs. As both model size and computational resources scale up, BLaST consistently outperforms the dense baseline. BLaST-95% achieves a speedup of 2.2x on 16 GPUs.
@@ 5.4. BLaST for Inference
Pretraining neural networks from scratch demands substantial computational resources. The widespread availability of pretrained LLMs can be leveraged rather than retrain a network from scratch. This section reports model accuracy across tasks under block sparsity at inference. We organize results by task and evaluate multiple models, datasets, and block sparsification algorithms. The goal is to show that BLaST is effective across a broad range of practical LLM applications.
5.4.1. Text Summarization
The DialogSum (Chen2021c, ) dataset tests whether an LLM can summarize a dialogue between two speakers. Table 1 shows the ROUGE-SUM(Lin2004, ) score when using Llama3.2-1B (Touvron2023a, ). and Qwen3-1.7B (Yang2025a, ) on DialogSum. Block size $1$ denotes fine-grained sparsity followed by block sparsity with block size ranging from 16 to 128.
The pruning algorithm influences the ROUGE-SUM scores achieved by the model for a given block size and target sparsity. oLLM, described in Section 3.3, improves scores for Llama3.2 and Qwen3. For the blocked prune and grow algorithm (Section 3.2), fine-grained sparsity (i.e. block size 1) shows the worst ROUGE-SUM scores. Increasing the block size from 16 to 128 improves the score. For the oLLM algorithm, the best ROUGE-SUM score is obtained for block size 1, whereas a larger block size shows deterioration in the ROUGE-SUM score.
We found that a procedure of dense finetuning followed by sparsification repeated multiple times works best for this dataset. This cycle of dense-sparse fine-tuning is done twice for all data points, leading to a total of 4 epochs of fine-tuning over the dataset. The dense baseline is computed using a single epoch. While accurate sparsification requires four times the fine-tuning cost of the dense model, the substantial inference speedups demonstrated in Section 5.4 justify this additional one-time investment, as inference constitutes the majority of LLM operational costs.
Model Pruning alg. $s_{max}$ Block size ($b$) Dense
1 16 32 64 128
Llama3.2
1B P&G(3.2) 70% 31.99 37.81 38.14 38.35 38.11 37.43
80% 31.55 36.28 37.47 37.41 37.15
90% 31.37 35.39 35.78 35.59 34.91
95% 31.15 34.5 34.54 33.84 33.52
oLLM (3.3) 70% 38.64 37.33 37.15 36.9 37.13
80% 37.83 36.58 36.43 36.4 36.7
90% 36.93 34.93 35.16 33.57 33.45
95% 35.89 33.15 33.5 31.85 31.68
Qwen3
1.7B P&G(3.2) 70% 30.73 38.68 38.89 38.35 38.28 41.22
80% 29.87 37.73 37.86 37.41 37.15
90% 30.53 36.48 35.78 35.59 34.49
95% 30.85 35.16 34.54 33.84 33.82
oLLM (3.3) 70% 40.82 38.95 39.05 38.97 38.85
80% 40.11 38.34 37.69 37.57 36.97
90% 38.75 36.68 35.93 35.73 35.92
95% 37.52 35.34 34.52 34.75 34.52
Table 1. ROUGE-SUM (Lin2004, ) score for text summarization for the DialogSum dataset using Llama3.2 and Qwen3. ROUGE-SUM measures the longest common subsequence (LCS) between a generated summary and reference summary, evaluating how well the generated summary captures the content and ordering of the reference.
This results can be attributed to the way in which these algorithms perform sparsification. The P&G is essentially a magnitude pruning algorithm that first prunes the weights with the least magnitude, and then grows back some of them using the heuristic of the weight gradient with large magnitudes. This heuristic provides only a rough approximation of the model’s curvature. Therefore, increasing the block size has the impact of incorporating weights that otherwise might have been removed when using fine-grained sparsity. On the other hand, oLLM uses second order curvature information to estimate weight importance more accurately. As a result, it can identify important weights with fine grained sparsity. Using 2D average pooling to aggregate sensitivity scores introduces inaccuracy by assuming that entire weight blocks can be pruned based on their average sensitivity. This approximation reduces the ROUGE-SUM score for oLLM under blocked sparsity.
5.4.2. Text Classification (GLUE)